### v1

# Efficient computation of stationary measures and the Lyapunov Landscape for families random dynamical systems with smooth additive noise

Abstract
We present an efficient and validated method for approximating the stationary measures of random dynamical systems with smooth additive noise. The approach leverages the strong regularizing properties of the associated transfer operator through a finite-dimensional reduction based on Fourier approximation. Explicit error bounds make the method suitable for use in computer-assisted proofs and rigorous numerical investigations; in particular, its efficiency enables systematic explorations of parameter space.
The method provides access to the stationary measure and supports the analysis of key statistical properties of the system. As an application, we study noise-induced phenomena, focusing on the transition from positive to negative Lyapunov exponent (commonly known as Noise Induced Order) in families of random unimodal maps with Gaussian additive noise.
By analyzing the Lyapunov exponent as a function of the system parameters, we identify transitions along a hypersurface in parameter space. The parameters we consider include the standard deviation (intensity) of the Gaussian noise and the shape of the unimodal map.
1. Introduction
In random dynamical systems, Noise Induced Order refers to the transition from positive to negative Lyapunov exponent as the intensity of the noise increases. This phenomenon was first observed numerically in 1983 in the Matsumoto–Tsuda model for the Belousov–Zhabotinsky chemical reaction [30]. This illustrates a nontrivial effect: the system appears chaotic for very small noise intensity, but as the noise increases, it displays increasingly regular behavior, eventually undergoing a transition to a negative Lyapunov exponent. Similar behavior has been observed for other indicators of chaos.
Despite the broad impact of these findings–with over 450 citations to [30] at the time of writing–a rigorous mathematical proof of such phenomena only appeared recently, using computer-assisted techniques [19]. Further developments include the demonstration of multiple transitions in the Lasota–Mackey map [12] and studies of similar behavior in unimodal maps with bounded variation noise [31], and in random skew-product systems over non-uniformly expanding bases [5]. Most recently, the phenomenon has also been investigated in high-dimensional systems with structured stochastic dynamics [11], indicating its robustness across different classes of random dynamical systems.
Alongside Noise Induced Order, other noise-induced phenomena such as Noise Induced Chaos have been studied in the applied literature [35].
Despite the number of studies on these phenomena, a rigorous mathematical understanding remains limited. The computer-assisted strategy developed in [19] has proven to be a powerful tool to investigate such questions. These phenomena are strongly tied to the statistical properties of typical trajectories–e.g., the sign of the Lyapunov exponent in one-dimensional maps depends on whether expansion or contraction dominates along typical orbits.
These statistical properties, in turn, are governed by the system’s stationary measures (as we discuss in Section 4), which can be characterized as fixed points of suitable transfer operators (see Section 5). The presence of noise has a regularizing effect on the transfer operators, making them and their fixed points approximable via finite-element methods: the core of the computer-assisted strategies we use. In this sense, noise helps the computer capture the dynamics at finite resolution in a way that enables rigorous inference about the underlying system.
Computer-assisted approaches to the approximation of invariant measures via discretization of the transfer operator have been extensively developed in the deterministic case, beginning with Ulam’s pioneering ideas [37]. See, e.g., [1, 4, 7, 13, 14, 20, 21, 22, 26, 27, 39] for developments using various approximation schemes, including Fourier and polynomial bases. Other strategies relying on periodic orbits have also been explored [32].
The strategy we apply here to get explicit estimates on the finite elements reduction error, relies on the a posteriori approach proposed in [21], allowing the convergence to equilibrium to be tested on the discretized transfer operator. With this approach the convergence to equilibrium estimate is performed by the computer instead of being given a priori by general considerations on the system. To allow such a complicated estimate to be performed efficiently we also implement a multi resolution approach, we called ”coarse-fine strategy” ([20],[23]) where the convergence of equilibrium of a large finite state markov chain is estimated by the convergence of equilibrium of a lower resolution version of it satisfying the same Lasota-Yorke-Doeblin-Fortet general estimates.
From an abstract computability perspective, [17] shows that even some computable deterministic systems may lack computable invariant measures. Similar noncomputability results are known for physical measures of smooth maps [33]. However, such barriers are removed in the presence of random noise: [8] shows that in a compact phace space, random systems with additive noise supported on a small ball always admit computable stationary measures.
Nevertheless, efficient algorithms for approximating stationary measures in the random setting remain relatively scarce [19, 29]. This paper aims to provide a highly efficient solution to this problem when the noise is Gaussian. While results are available for stochastic differential equations (see [10]), and in the SDE case alternatives such as adjoint methods exist [9], our approach is based directly on discretization of the transfer operator using a Fourier scheme.
We demonstrate the effectiveness of the method with high-accuracy enclosures of the stationary measure and Lyapunov exponent in concrete one-dimensional examples. We then apply these tools to study transitions such as Noise Induced Order and Noise Induced Chaos.
2. Summary of the Results
This paper has two main contributions:
  • We develop a general theory for approximating the stationary measure of random dynamical systems with Gaussian noise in $L^{2}([-1,1])$, including rigorous error bounds based on a Fourier finite-element discretization of the transfer operator. This framework allows one to rigorously enclose the Lyapunov exponent.
  • We apply this method to study a family of unimodal maps previously introduced in [31], demonstrating the presence of multiple noise-induced transitions and visualizing the landscape of the Lyapunov exponent in parameter space.
The following computability result underpins the general theory:
Main result 1.
Let $T:[-1,1]\to\mathbb{R}$ be a measurable map such that the pushforward of Lebesgue measure by $T$ is absolutely continuous with respect to Lebesgue measure (i.e., $T$ is non-singular). Define the random dynamical system with Gaussian noise:
$X_{n+1}=\tau(T(X_{n})+\Omega_{n}),$
where $\Omega_{n}\sim\mathcal{N}(0,\sigma^{2})$ are i.i.d. random variables, and $\tau:\mathbb{R}\to[-1,1]$ is a periodic boundary condition.
Then, for every $\sigma>0$, the system admits a unique stationary measure $\mu_{\sigma}=f_{\sigma}m$, with analytic density $f_{\sigma}$, where $m$ is the Lebesgue measure on [1,1][-1,1].
Moreover, for every observable $\phi\in L^{2}(m)$, the Birkhoff average
$\frac{1}{N}\sum_{n=0}^{N-1}\phi(X_{n})$
converges almost surely to $\int\phi\,d\mu_{\sigma}$, and this integral can be rigorously enclosed with arbitrary precision using the algorithm presented in Section 8.
The proof of this result is developed throughout the paper. Existence of the stationary measure is established in Lemma 6.3, while uniqueness is shown in Proposition 6.8 as a consequence of the Doeblin argument used to estimate mixing rates in Proposition 6.6. Analyticity of the stationary density is proved in Corollary 7.9. Furthermore, the same Doeblin-type estimates yield the computability results discussed in Section 9, namely, the computability of the stationary measure in Theorem 9.4 and of Birkhoff averages in Corollary 9.5.
As an application, we consider a discrete-time random dynamical system defined by the two-parameter family of maps
(1) $T_{\alpha,\beta}(x)=\beta-(1+\beta)|x|^{\alpha}$
with Gaussian additive noise. Some sufficient conditions for a transition of Noise Induced Order type were studied in [31] in some case of bounded variation noise. Here in the Gaussian case we provide a more refined description of both Noise Induced Order and Noise Induced Chaos exploring a whole region of the parameter space.
We rigorously enclose regions of parameter space where the Lyapunov exponent transitions from positive to negative, proving the existence of multiple noise-induced transitions.
This leads to the following computer-assisted result:
Main result 2.
Let $T_{\alpha,\beta}:[-1,1]\to\mathbb{R}$ be as above, and let $\tau:\mathbb{R}\to[-1,1]$ denote the projection modulo 2. Consider the random system
$X_{n+1}=\tau(T_{\alpha,\beta}(X_{n})+\Omega_{n}),$
where $\Omega_{n}\sim\mathcal{N}(0,\sigma^{2})$ are i.i.d. Gaussian variables.
Then the top Lyapunov exponent $\lambda(\alpha,\beta,\sigma)$ is well-defined for all $\alpha\in[1,\infty),\beta\in(-1,1],\sigma>0$. Moreover:
  • For $\beta=1$, the function $\alpha\mapsto\lambda(\alpha,1,\sigma)$ changes sign as $\sigma$ increases, for $\alpha\in[3,4]$, from positive to negative.
  • For $\alpha=3$, the function $\beta\mapsto\lambda(3,\beta,\sigma)$ exhibits multiple sign changes as $\sigma$ increases, for $\beta\in[0.8447,0.8694]$.
Remark 2.1.
The proof of Main Result 2 is obtained by rigorously enclosing the Lyapunov exponent using validated numerics.
Specifically:
  • For $\beta=1$, the Lyapunov exponent was enclosed rigorously at $\alpha=3+\frac{i}{1024}$, for $i=0,\ldots,1024$, $\sigma=1/16+15/16\frac{j}{1024}$ for $j=0,\ldots,1024$.
  • For $\alpha=3$, the Lyapunov exponent was enclosed rigorously at $\beta=\frac{51}{64}+\frac{i}{8192}$, for $i=1,\ldots,1024$, $\sigma=1/16+15/16\frac{j}{1024}$ for $j=0,\ldots,1024$.
It is known that in the present case, considering smooth families of systems with additive smooth noise and periodic boundary conditions ([24]) the stationary measure and hence the Lyapunov exponent vary in a smooth way as the parameters vary. The following figures show in the case of the family 1 with Gaussian noise, how the sign of the Lyapunov exponent varies with the parameters, by rigorously computed enclosures produced by our code:
  • Figure 1 shows the regions in the $(\alpha,\sigma)$-plane where the Lyapunov exponent transitions from positive (red) to negative (blue).
  • Figure 2 shows the analogous transition in the $(\beta,\sigma)$-plane for fixed $\alpha=3$.
Figure 1. Parameter region in the $(\alpha,\sigma)$-plane. Blue: $\lambda<0$, Red: $\lambda>0$. Rigorous enclosures for the Lyapunov exponent at the points marked with black crosses can be found in Table 1. The Lyapunov exponent was rigorously enclosed on the points of the grid $\alpha=3+\frac{i}{1024}$, for $i=0,\ldots,1024$, $\sigma=1/16+15/16\frac{j}{1024}$ for $j=0,\ldots,1024$
Figure 2. Parameter region in the $(\beta,\sigma)$-plane for $\alpha=3$. Blue: $\lambda<0$, Red: $\lambda>0$. Rigorous enclosures for the Lyapunov exponent at the points marked with black crosses can be found in Table 2. The Lyapunov exponent was enclosed rigorously on the points of the grid $\beta=\frac{51}{64}+\frac{i}{8192}$, for $i=1,\ldots,1024$, $\sigma=1/16+15/16\frac{j}{1024}$ for $j=0,\ldots,1024$.
We emphasize that both Figure 1 and Figure 2 are obtained using fully rigorous numerical methods. As such, up to errors introduced by graphical rendering (e.g., rasterization or color interpolation), they can be interpreted as certified results. In particular, they provide an effective visualization of Main Result 2 stated above. The crosses in the figures correspond to sample enclosures of the Lyapunov exponent presented in Tables 1 and 2, we report as examples of the computation done, presenting explicit values and enclosure bounds; the data is available at [18].
The values reported in Table 1 were computed to machine precision using a Galerkin truncation at frequency $128$, with a total runtime of $2.02$ seconds on a single thread of a Ryzen 5 5600 processor. No parallelization was used, and the annealed transfer operator was recomputed independently for each pair of parameter values. Consequently, the average runtime for the full computation -including operator discretization, mixing time estimation, and rigorous Lyapunov exponent evaluation- is approximately $0.5$ seconds per parameter point. Despite its simplicity, this naive implementation already delivers high performance, achieving $15$ certified digits without requiring any specialized optimization.
For the full parameter space plots (Figures 1 and 2), we employed only minimal optimization: in particular, we computed the transfer operator of the deterministic map once for each fixed value of $\alpha$ or $\beta$, and reused it across all $\sigma$ values. This reduced computational cost while preserving the rigor of the enclosures. As a result, our algorithm remains both validated and efficient, and is well suited for the systematic exploration of large parameter regions in a robust way.
We stress again that both the algorithm and the discretization scheme are independent of the specific dynamical properties of the underlying deterministic map. This reflects an instance of the a posteriori validation philosophy, which in our context evolves from the coarse-fine strategy developed in [20]. The essential analytical estimates—such as mixing and spectral contraction—are established through rigorous linear algebra computations performed on the discretized operator. As a result, our method is problem-agnostic within the class of random dynamical systems with Gaussian noise superimposed on non-singular interval maps. This structural independence underlines the generality and flexibility of our approach.
Label $\alpha$ $\sigma$ $\lambda$
$x_{1}$ $3.25$ $0.2$ $0.139610862369467\pm 10^{-15}$
$x_{2}$ $3.5$ $0.2$ $0.047682027067898\pm 10^{-15}$
$x_{3}$ $3.5$ $0.4$ $0.095573727164159\pm 10^{-15}$
$x_{4}$ $3.8$ $0.4$ $0.223066357002470\pm 10^{-15}$
Table 1. Sample values in Figure 1
Label $\beta$ $\sigma$ $\lambda$
$x_{1}$ $0.875$ $0.2$ $0.058040752469902\pm 10^{-15}$
$x_{2}$ $0.9$ $0.2$ $0.098888825866413\pm 10^{-15}$
$x_{3}$ $0.875$ $0.4$ $0.036288212585984\pm 10^{-15}$
$x_{4}$ $0.9$ $0.4$ $0.003180843719330\pm 10^{-15}$
Table 2. Sample values in Figure 2
It is worth emphasizing that linear response methods have been developed and adapted to the class of perturbations studied in this paper [16, 24]. In particular, the theory of linear response for systems perturbed by Gaussian noise, as developed in [24], combined with the implicit function theorem, supports our claim that the zeros of the Lyapunov exponent lie along a hypersurface in parameter space. A more detailed investigation of the dynamics associated with parameters on this hypersurface presents a promising direction for future research, as such systems may exhibit intermittent-like behavior.
Indeed, recent work by Goverse and collaborators [25] shows that, at the onset of chaos (i.e., when the Lyapunov exponent vanishes) in random circle maps, one observes intermittent two-point dynamics and infinite invariant measures. Intermittent two point behavior is exactly the behavior that may arise in our setting along the hypersurface where the Lyapunov exponent is zero. Numerical evidence of similar intermittent two-point dynamics has also been observed in our simulations; see Figure 3. In Table 3, we report validated enclosures for values of the noise amplitude $\sigma$ as a function of the shape parameter $\alpha$ at which this transition occurs, and in Figure 4, we display an enclosure of the function $\sigma(\alpha)$ at sample points. We stress, however, that our current computations only establish the existence of such transitions, without addressing their uniqueness or potential multiplicity.
Figure 3. Numerical experiments plotting $|T^{i}_{\omega}x_{0}-T^{i}_{\omega}(y_{0})|$ for initial conditions $x_{0}=-0.1,y_{0}=0.9$ for noise realizations with $\sigma=0.1,0.2971290840221,0.4$; the transition from positive to negative Lyapunov exponent was rigorously enclosed in $[0.2971290840221,0.2971290840222]$.
Figure 4. Enclosure of the transition value $\sigma$ as a function of $\alpha$; in blue the lower bound, in red the upper bound. This figure gives a graphical representation of the definition at which the hypersurface of zero Lyapunov exponent is known in our rigorous data.
Table 3. Sample of detected Lyapunov Crossings
Row $\alpha$ $\sigma_{1}$ $\lambda_{1}$ $\sigma_{2}$ $\lambda_{2}$
11 $3.0$ $0.523926$ $0.00107101,0.00107102]_{256}$ $0.525757$ $0.0005515,-0.000551499]_{256}$
22 $3.04883$ $0.498291$ $0.000820052,0.000820053]_{256}$ $0.500122$ $0.000905828,-0.000905827]_{256}$
33 $3.09766$ $0.474487$ $0.000816379,0.00081638]_{256}$ $0.476318$ $0.000986452,-0.000986451]_{256}$
44 $3.14648$ $0.452515$ $0.000518536,0.000518537]_{256}$ $0.454346$ $0.00133441,-0.0013344]_{256}$
55 $3.19531$ $0.430542$ $0.00132734,0.00132735]_{256}$ $0.432373$ $0.000545068,-0.000545067]_{256}$
66 $3.24414$ $0.4104$ $0.000980779,0.00098078]_{256}$ $0.412231$ $0.000883612,-0.000883611]_{256}$
77 $3.29297$ $0.390259$ $0.000958772,0.000958773]_{256}$ $0.39209$ $0.000864106,-0.000864105]_{256}$
88 $3.3418$ $0.370117$ $0.000848431,0.000848432]_{256}$ $0.371948$ $0.000897693,-0.000897692]_{256}$
99 $3.39062$ $0.349976$ $0.000214735,0.000214736]_{256}$ $0.351807$ $0.00141992,-0.00141991]_{256}$
$10$ $3.43945$ $0.328003$ $0.000105052,0.000105053]_{256}$ $0.329834$ $0.00136926,-0.00136925]_{256}$
$11$ $3.48828$ $0.302368$ $0.000835062,0.000835063]_{256}$ $0.304199$ $0.000416835,-0.000416834]_{256}$
$12$ $3.53711$ $0.273071$ $0.000731332,0.000731333]_{256}$ $0.274902$ $0.000262686,-0.000262685]_{256}$
$13$ $3.58594$ $0.23645$ $0.00018605,0.000186051]_{256}$ $0.238281$ $0.000581865,-0.000581864]_{256}$
$14$ $3.63477$ $0.194336$ $0.000300537,0.000300538]_{256}$ $0.196167$ $0.000607178,-0.000607177]_{256}$
$15$ $3.68359$ $0.165039$ $0.000930034,0.000930035]_{256}$ $0.16687$ $0.000639231,-0.00063923]_{256}$
$16$ $3.73242$ $0.146729$ $0.00171531,0.00171532]_{256}$ $0.14856$ $0.000694382,-0.000694381]_{256}$
$17$ $3.78125$ $0.133911$ $0.00236626,0.00236627]_{256}$ $0.135742$ $0.000913736,-0.000913735]_{256}$
$18$ $3.83008$ $0.124756$ $0.00134207,0.00134208]_{256}$ $0.126587$ $0.00273638,-0.00273637]_{256}$
1919 $3.87891$ $0.115601$ $0.00469925,0.00469926]_{256}$ $0.117432$ $0.000340576,-0.000340575]_{256}$
2020 $3.92773$ $0.110107$ $0.00127257,0.00127258]_{256}$ $0.111938$ $0.00444233,-0.00444232]_{256}$
2121 $3.97656$ $0.104614$ $1.11679\times 10^{-5},1.1168\times 10^{-5}]_{256}$ $0.106445$ $0.00644444,-0.00644443]_{256}$
The full results of the computation are accessible at Harvard Dataverse [18]; the data is provided both as a dataframe in JLD2 format and in CSV format. The $l^{2}$ mixing rates for the iterations of the discretized operator are provided, which can be used to estimate the mixing rates for the abstract annealed operator or as an ingredient in further rigorous computations, as those of linear response and diffusion coefficients. In figure 5 we present heatmaps built upon the rigorous mixing bounds in the $\alpha$, $\sigma$ plane when $\beta=1.0$.
Figure 5. Upper bounds on $\log(||P^{N}_{k}|_{U_{0}}||_{2})$ for $N=2$ and $N=5$ as a function of $\alpha$ and $\sigma$.
Beyond the specific examples considered in this work, our approach contributes to the broader understanding of random dynamical systems exhibiting nonuniformly hyperbolic behavior. By combining transfer operator techniques, Fourier-based discretization, and validated numerics, we are able to rigorously analyze statistical properties and dynamical transitions even in the absence of uniform expansion.
Future directions include the study of finer statistical features, such as variance, limit theorems, and response functions, following the lines of [2, 3], as well as a detailed spectral analysis of the associated transfer operators. The latter can be approached using the pseudospectral and rigorous numerical techniques recently developed in [6].
3. Logistic Maps with noise
In this section we define more precisely the systems we are going to consider. In this work, we consider the following family of maps, generalization of the quadratic family:
(2) $T_{\alpha,\beta}(x)=\beta-(1+\beta)|x|^{\alpha},$
where $\alpha\geq 1$, $\beta\in(-1,1]$ and $x\in[-1,1]$.
We introduce the notations we use for the Gaussian distribution, which will be necessary to define the random dynamical system.
Definition 3.1.
A random variable $\Omega$ is said to have a standard Gaussian distribution if its density function is given by
$\rho_{\sigma}(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-x^{2}/(2\sigma^{2})}.$
While the Gaussian distribution has unbounded support, the empirical rule (also known as the 68–95–99.7 rule) indicates that the standard deviation effectively quantifies the typical size of noise fluctuations.
Definition 3.2.
Define $\tau:\mathbb{R}\rightarrow[-1,1)$ by
$\tau(x)=x-2\left\lfloor\frac{x+1}{2}\right\rfloor.$
We refer to $\tau$ as the periodic boundary condition.
Remark 3.3.
The function $\tau$ maps $\mathbb{R}$ onto [1,1)[-1,1) by identifying points modulo the equivalence relation
$x\sim y\quad\text{if }x-y=2k,\quad k\in\mathbb{Z}.$
With these definitions in place, we can now describe the dynamical system with Gaussian noise
(3) $X_{n+1}=\tau(T_{\alpha,\beta}(X_{n})+\Omega_{\sigma}(n)),$
where $\Omega_{\sigma}(n)$ are random variables i.i.d. with Gaussian distribution and $\tau$ the boundary condition.
4. Lyapunov exponent for the r.d.s.
In this section, we recall basic definitions for random dynamical systems with additive noise on the interval [1,1][-1,1], including the Lyapunov exponent.
We model the system as a skew product:
$F:\mathbb{R}^{\mathbb{N}}\times[-1,1]\rightarrow\mathbb{R}^{\mathbb{N}}\times[-1,1],\quad(\omega,x)\mapsto\left(s(\omega),\tau(T_{\alpha,\beta}(x)+\omega_{0})\right),$
where $T_{\alpha,\beta}:[-1,1]\to\mathbb{R}$ is a measurable map, $\tau:\mathbb{R}\to[-1,1]$ denotes the periodic projection (e.g., modulo $2$), and $s:\mathbb{R}^{\mathbb{N}}\to\mathbb{R}^{\mathbb{N}}$ is the left shift: $s(\omega)_{n}=\omega_{n+1}$.
The space $\Omega:=\mathbb{R}^{\mathbb{N}}$, representing the realizations of the noise process, is endowed with the product topology and the corresponding Borel $\sigma$-algebra. We define the product probability measure $\nu:=\rho_{\sigma}^{\mathbb{N}}$, where $\rho_{\sigma}$ is the Gaussian measure $\mathcal{N}(0,\sigma^{2})$ on $\mathbb{R}$. We refer to $(\Omega,\nu)$ as the noise space.
The random dynamical system is then described by the iterates of $F$, acting on the product space $\Omega\times[-1,1]$, with randomness encoded in the noise sequence $\omega\in\Omega$.
We are interested in measuring the average expansion along orbits of the skew product, that is, in studying the limit
$\lim_{n\to\infty}\frac{1}{n}\log\left|\frac{\partial F^{n}}{\partial x}(\omega,x)\right|.$
We note that the function $\tau$ is piecewise linear with jump discontinuities at the points x=2k1x=2k-1, $k\in\mathbb{Z}$, where the floor function in its definition causes discontinuous jumps. Consequently, $\tau$ is not differentiable at these points. However, since the set of discontinuities is countable, it has Lebesgue measure zero. Therefore, the chain rule
$\frac{d}{dx}\tau(T_{\alpha,\beta}(x)+y)=T^{\prime}_{\alpha,\beta}(x)$
holds for almost every $x$, provided that $T_{\alpha,\beta}$ is differentiable and the argument $T_{\alpha,\beta}(x)+y$ avoids the discontinuity set of $\tau$ for almost all $y$. Since the noise distribution is absolutely continuous, this is indeed the case for almost every realization.
Thus, letting $\mathrm{proj}_{x}$ denote the projection onto the second coordinate, and setting $X_{i}:=\mathrm{proj}_{x}(F^{i}(\omega,x))$, we find
$\frac{1}{n}\log\left|\frac{\partial F^{n}}{\partial x}(\omega,x)\right|=\frac{1}{n}\sum_{i=0}^{n-1}\log|T^{\prime}_{\alpha,\beta}(X_{i})|.$
Definition 4.1.
A probability measure $\mu$ on [1,1][-1,1] is called stationary if the product measure $\nu\times\mu$ is invariant under the skew product $F$.
Definition 4.2.
A stationary measure $\mu$ is said to be ergodic if $\nu\times\mu$ is ergodic with respect to the skew product $F$.
Lemma 4.3.
Suppose the random dynamical system admits a unique stationary measure $\mu_{\sigma}$. Let $\phi:\mathbb{R}^{\mathbb{N}}\times[-1,1]\to\mathbb{R}$ be a measurable function such that $\phi(\omega,x)$ depends only on $x$, i.e., there exists $\bar{\phi}$ such that $\phi(\omega,x)=\bar{\phi}(x)$. Suppose $\bar{\phi}\in L^{1}(\mu_{\sigma})$, then for $\mu_{\sigma}$-almost every $x_{0}$ and for $\nu$-almost every $\omega$,
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}\phi(F^{i}(\omega,x_{0}))=\int\bar{\phi}\,d\mu_{\sigma}.$
Proof.
The ergodic decomposition theorem for stationary measures [38, Theorem 5.14] asserts that every stationary measure is a convex combination of ergodic stationary measures. Since $\mu_{\sigma}$ is unique, it must be ergodic. Because $\bar{\phi}\in L^{1}(\mu_{\sigma})$, it follows that $\phi\in L^{1}(\nu\times\mu_{\sigma})$. Therefore, by the Birkhoff Ergodic Theorem,
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}\phi(F^{i}(\omega,x_{0}))=\int\phi(\omega,x)\,d(\nu\times\mu_{\sigma})=\int\bar{\phi}(x)\,d\mu_{\sigma}$
for $(\omega,x_{0})$ in a set of full $\nu\times\mu_{\sigma}$-measure. ∎
In our setting, the function $\phi(\omega,x)=\log|T^{\prime}_{\alpha,\beta}(x)|$ depends only on the second coordinate and satisfies the integrability condition required by Lemma 4.3. Therefore, the Birkhoff average converges almost surely to the space average with respect to $\mu_{\sigma}$, and we obtain the following characterization of the Lyapunov exponent:
(4) $\lim_{n\to\infty}\frac{1}{n}\log\left|\frac{\partial F^{n}}{\partial x}(\omega,x)\right|=\int\log|T^{\prime}_{\alpha,\beta}|\,d\mu_{\sigma}.$
Thus, when the stationary measure is unique, the Lyapunov exponent can be computed by integrating the logarithm of the derivative of the map with respect to the stationary measure.
As $\sigma$ varies, we are interested in how the Lyapunov exponent behaves as a function of noise amplitude and its dependence on the parameters $\alpha$ and $\beta$, i.e., studying the following function
$\lambda(\alpha,\beta,\sigma)=\int_{-1}^{1}\log|T^{\prime}_{\alpha,\beta}|\,d\mu_{\sigma};$
the next sections will be devoted to proving it is well-defined and some of its properties.
Definition 4.4.
We say that a system exhibits a transition of the Lyapunov exponent if there exist $0<\sigma_{1}<\sigma_{2}$ such that for all $\sigma\geq\sigma_{1}$, the system admits a unique stationary measure with density $f_{\sigma}$, and the Lyapunov exponent transitions continuously changing sign, that is, $\lambda(\sigma_{1})\cdot\lambda(\sigma_{2})<0$. We say the system exhibits a multiple transition if several such sign transitions occur.
5. The annealed transfer operator
Following a well-established strategy, we want to reframe the problem of finding a stationary measure as a fixed point problem for an appropriate operator acting on measures, the annealed transfer operator In this section we define the operator and show that is regularizing, i.e., preserves a space of regular densities. We start by investigating the properties of convolution with a Gaussian kernel.
Definition 5.1.
Denote by $\chi_{X}$ the characteristic function of the set $X$. In the following we will denote by
$\widehat{f}(x)=f(x)\cdot\chi_{[-1,1]}(x)$
the function that extend $f$ by 0 outside its intervals of definition. Given a probability measure $\mu$ on [1,1][-1,1], we define its extension $\widehat{\mu}$ on $\mathbb{R}$ as the unique measure $\widehat{\mu}$ on $\mathbb{R}$ such that
$\widehat{\mu}(A)=\mu(A\cap[-1,1])$
for all $A$ measurable in $\mathbb{R}$.
We will define an operator as the convolution of a measure with $\rho_{\sigma}$, and we will examine some of its regularizing properties. The most important of these is Proposition 5.4, which tells us that the convolution of a measure with $\rho_{\sigma}$ is absolutely continuous with respect to the Lebesgue measure.
Definition 5.2.
Let $\mu$ be any probability measure on [1,1][-1,1], and $\rho_{\sigma}$ the Gaussian distribution; their convolution $\rho_{\sigma}*\widehat{\mu}$ is the unique probability measure on $\mathbb{R}$ such that
$\rho_{\sigma}*\widehat{\mu}(A)=\int_{-\infty}^{+\infty}\rho_{\sigma}(y)\widehat{\mu}(A-y)dm(y),$
where $A-y:=\{x-y|x\in A\}$.
The following lemma, provided without proof, states some properties of extension and convolution that follow directly from the definition.
Lemma 5.3.
Let $\mu$ be a probability measure on [1,1][-1,1]
  1. (1)
    if $\mu=fdm$ we have that $\widehat{\mu}=\widehat{f}dm$,
  2. (2)
    $\rho_{\sigma}*\widehat{\mu}(\mathbb{R})=1$,
  3. (3)
    if $\mu=fdm$ then $\rho_{\sigma}*\widehat{\mu}$ has density $\rho_{\sigma}*\widehat{f}$ (i.e. $\rho_{\sigma}*\widehat{\mu}=(\rho_{\sigma}*\widehat{f})dm$).
The proofs of these lemmas follow immediately from the definition. We show now a simple result, proving that the convolution with a Gaussian kernel maps measures into the more regular space of absolutely continuous invariant measures.
Proposition 5.4.
Let $\mu$ be a probability measure on [1,1][-1,1], then $\rho_{\sigma}*\widehat{\mu}$ is a probability measure on $\mathbb{R}$, absolutely continuous with respect to Lebesgue.
Proof.
Let $\Phi:L^{1}(m)\rightarrow\mathbb{R}$ defined by
[…]
where $h\in L^{1}(m)$.
[…]
i.e., $\Phi$ is a bounded linear functional on $L^{1}(m)$.
Then from [34, Theorem 6.16] there is a unique $g\in L^{\infty}(m)$, such that
$\int_{\mathbb{R}}hd(\rho_{\sigma}*\widehat{\mu})=\int_{\mathbb{R}}hgdm$
for all $h\in L^{1}(m)$, i.e., $\rho_{\sigma}*\widehat{\mu}=gdm$. ∎
Definition 5.5.
Let $\mathcal{B}([-1,1])$ be the Borel $\sigma$-algebra on [1,1][-1,1], defined as the restriction of the Borel $\sigma$-algebra of $\mathbb{R}$ to [1,1][-1,1].
Let $\mu$ be a measure on $([-1,1],\mathcal{B}([-1,1]))$, a measurable map $T:[-1,1]\rightarrow[-1,1]$ is called non-singular with respect to $\mu$ if, $\mu(T^{-1}(A))=0$ for all $A\in\mathcal{B}([-1,1])$ such that $\mu(A)=0$.
We define the Annealed Transfer Operator associated to the system with noise (see e.g. [28] or [15] for an introduction to the topic).
Definition 5.6.
Let $T:[-1,1]\rightarrow[-1,1]$ be a measurable map. The map $T$ induces an operator on $L:\mathcal{SM}([-1,1])\rightarrow\mathcal{SM}([-1,1])$ where $\mathcal{SM}([-1,1])$ is the space of signed measures on [1,1][-1,1], defined in the following way: if $\mu\in\mathcal{SM}([-1,1])$ then
$L\mu(A)=\mu(T^{-1}A)$
for all measurable set $A$. This operator is called the transfer operator associated to $T$.
The space of Lebesgue absolutely continuous measures is a vector subspace of
$\mathcal{SM}([-1,1])$; if $T$ is non-singular with respect to Lebesgue then $L$ preserves this subspace of absolutely continuous measures and induces an operator from $L^{1}([-1,1])$ into itself called the Perron-Frobenius operator. We will denote by $P$ the Perron-Frobenius operator.
Remark that, even though the dynamical system $T$ is defined on the interval [1,1][-1,1], the addition of random noise may result in iterates lying outside this interval. The next definition fixes the notation for the action of the boundary condition on measures.
Definition 5.7.
Let $\tau:\mathbb{R}\rightarrow[-1,1]$ be the boundary condition as in the Definition 3.2. We denote by $\tau_{*}$ the push-forward map acting on measures by
$(\tau_{*}\mu)(A)=\mu(\tau^{-1}(A)),$
i.e., the transfer operator associated to the boundary condition.
Definition 5.8.
The annealed transfer operator $L_{\sigma}$ associated to the system with noise is defined by
[…]
The next corollary follows directly from Lemma 5.4 and the definitions.
Corollary 5.9.
The operator $L_{\sigma}$ induces an operator $P_{\sigma}:L^{1}(m)\to L^{1}(m)$, acting on densities such that
[…]
We refer to this operator as the annealed Perron-Frobenius operator.
Proof.
Let $\mu=fdm$ be a absolutely continuous probability measure with density $f$.
By Definition 5.6 we have that
$L\mu=Pfdm.$
By Lemma 5.3, we have that
$\rho_{\sigma}*\widehat{L\mu}=(\rho_{\sigma}*\widehat{Pf})dm,$
where the Lebesgue measure on the right-hand side is defined on $\mathbb{R}$. Therefore, we have that
$\tau_{*}(\rho_{\sigma}*\widehat{L\mu})=\tau_{*}(\rho_{\sigma}*\widehat{Pf})dm,$
[…]
The following result connects the annealed transfer operator with the skew-product representation of a random dynamical system [38, Section 5.3].
Lemma 5.10.
Let $T$ be any non-singular dynamical map of the interval and let the random system with Gaussian noise be defined as:
(5) $X_{n+1}=\tau(T(X_{n})+\Omega_{\sigma}(n)),$
where $\Omega_{\sigma}(n)$ are random variables i.i.d. with Gaussian distribution with standard deviation $\sigma$ and $\tau$ are periodic boundary conditions. If $\mu_{\sigma}$ is a fixed point for $L_{\sigma}$, then $\mu_{\sigma}$ is a stationary measure for the random dynamical system.
Remark 5.11.
If $P_{\sigma}$ is the Perron-Frobenius operator operating on densities, and $f_{\sigma}$ is a fixed point of this operator
$P_{\sigma}f_{\sigma}=f_{\sigma}$
then $\mu_{\sigma}=f_{\sigma}dm$, where $m$ is the Lebesgue measures is a stationary measure.
The discussion in subsection 3 can now be restated in the following theorem; we refer also to [38, Proposition 5.4].
Theorem 5.12 (Birkhoff Ergodic Theorem for random dynamical system).
Suppose that $L_{\sigma}$ has a unique stationary measure $\mu_{\sigma}$ and let $\phi\in L^{1}(\mu_{\sigma})$. Then, for $\mu_{\sigma}$ almost every initial condition $X_{0}=x_{0}$ and with probability 11
$\displaystyle\lim_{n\rightarrow+\infty}\frac{1}{n}\sum_{i=0}^{n-1}\phi(X_{i})=\int\phi d\mu_{\sigma}.$
6. Regularization properties of the convolution
In section 5, we defined the operator $P_{\sigma}$. Our goal is to approximate this operator by a finite rank operator and use a matrix associated with this finite rank approximation to enclose a fixed point (section 8). To achieve this, we will introduce and use the Fourier transform, its truncation and examine some important properties of the finite rank matrix obtained by projecting through truncation (Proposition 7.6, Corollary 7.8), often called the Galerkin approximation.
Let $A$, $B$ be normed vector spaces and let $Q:A\rightarrow B$ be a linear operator.
We denote the mixed norm $\|Q\|_{A\rightarrow B}$ as
$\|Q\|_{A\rightarrow B}:=\sup_{f\in A,\|f\|_{A}\leq 1}\|Qf\|_{B}.$
Definition 6.1.
Define $N_{\sigma}:L^{1}\rightarrow L^{1}$ such that
$N_{\sigma}f(x)=(\tau_{*}(\rho_{\sigma}*\widehat{f}))(x).$
We will see that the operator $N_{\sigma}$ relates the Perron-Frobenius operator $P$ of the dynamical system $T$ with the operator $P_{\sigma}$, defined in Lemma 5.9, acting on densities.
Let $f\in L^{1}([-1,1])$. We have that
$N_{\sigma}f(x)$ $=(\rho_{\sigma}*\widehat{f})(\tau^{-1}(x))$
(6) $=\displaystyle\sum_{j=-\infty}^{+\infty}\int_{-1}^{1}\rho_{\sigma}(x+2j-y)f(y)dy.$
Note that, if $f\in L^{1}([-1,1])$, then
$N_{\sigma}Pf(x)=\tau_{*}((\rho_{\sigma}*\widehat{Pf})(x))=P_{\sigma}f(x)$
and therefore $N_{\sigma}P=P_{\sigma}$.
We summarize some properties of the operator $N_{\sigma}$ in the next lemma.
Lemma 6.2.
Let $\rho_{\sigma}$ be the Gaussian density with mean 0 and standard deviation $\sigma$ and $\tau$ be the periodic boundary condition and let
$),N_{\sigma}f=(\tau_{*}(\rho_{\sigma}*\widehat{f}$)),
for each function $f\in L^{1}(m)$ Then
  1. (1)
    $N_{\sigma}f(x)=\int_{-1}^{1}(\tau_{*}\rho_{\sigma})(x-y)f(y)dy$,
  2. (2)
    $N_{\sigma}$ is a bounded operator from $L^{1}\to L^{\infty}$,
  3. (3)
    $N_{\sigma}||_{L^{1}}\leq 1$,
  4. (4)
    $(N_{\sigma}f)^{\prime}(x)=\int_{-1}^{1}\tau_{*}(\rho^{\prime}_{\sigma})(x-y)f(y)dy$,
  5. (5)
    $N_{\sigma}$ is a bounded operator from $L^{1}\to BV([-1,1])$,
  6. (6)
    If $f\geq 0$, $\int_{-1}^{1}f=1$, then $N_{\sigma}f(x)>0$ for all $x\in[-1,1]$, and
    $N_{\sigma}f(x)\geq\frac{1}{\sigma\sqrt{2\pi}}e^{-1/(2\sigma^{2})}\quad\text{for all }x\in[-1,1].$
Proof.
We begin by analyzing $\tau_{*}\rho_{\sigma}(x)=\sum_{j=-\infty}^{+\infty}\rho_{\sigma}(x+2j):[-1,1]\to\mathbb{R}$. We define the sequence of partial sums:
$F_{N}(x):=\sum_{j=-N}^{N}\rho_{\sigma}(x+2j),\quad x\in[-1,1].$
Each $F_{N}$ is smooth, and
$F_{N}^{\prime}(x)=\sum_{j=-N}^{N}\rho_{\sigma}^{\prime}(x+2j).$
Since the Gaussian derivative satisfies
$|\rho_{\sigma}^{\prime}(x)|\leq\frac{|x|}{\sigma^{3}\sqrt{2\pi}}e^{-x^{2}/(2\sigma^{2})},$
we have uniform convergence of the series $\sum_{j\in\mathbb{Z}}\rho_{\sigma}(x+2j)$ and $\sum_{j\in\mathbb{Z}}\rho_{\sigma}^{\prime}(x+2j)$ on [1,1][-1,1]. It follows from the Weierstrass $M$-test that $F_{N}\to\tau_{*}\rho_{\sigma}$ and $F_{N}^{\prime}\to\tau_{*}\rho_{\sigma}^{\prime}$ uniformly.
Hence, $\tau_{*}\rho_{\sigma}$ is differentiable and:
$(\tau_{*}\rho_{\sigma})^{\prime}(x)=\sum_{j\in\mathbb{Z}}\rho_{\sigma}^{\prime}(x+2j)=\tau_{*}(\rho_{\sigma}^{\prime})(x),$
as claimed.
Then, by definition
$(\tau_{*}(\rho_{\sigma}*\widehat{f}))(x)$ $=(\rho_{\sigma}*\widehat{f})(\tau^{-1}(x))$
$=\displaystyle\sum_{j=-\infty}^{+\infty}(\rho_{\sigma}*\widehat{f})(x+2j)$
$=\displaystyle\sum_{j=-\infty}^{+\infty}\int_{-\infty}^{+\infty}\rho_{\sigma}(x+2j-y)f(y)\chi_{[-1,1]}(y)dy$
$=\int_{-1}^{1}\displaystyle\sum_{j=-\infty}^{+\infty}\rho_{\sigma}(x-y+2j)f(y)dy$
$=\int_{-1}^{1}(\tau_{*}\rho_{\sigma})(x-y)f(y)dy.$
We justify the interchange of the integral and the summation by Fubini-Tonelli, using the uniform convergence of the $F_{N}$.
[…]
Fix $g\in L^{1}$, $g\geq 0$, then
$\int_{-1}^{1}|N_{\sigma}g|dx=\int_{-1}^{1}\int_{-1}^{1}\tau_{*}\rho_{\sigma}(x-y)g(y)dxdy=\int g(y)dy,$
where the exchange in order of integrations is again justified by the uniform convergence of the $F_{N}$ and using the fact that $\int\rho_{\sigma}(x-y)dx=1$ for all $y$. Observing that the operator $N_{\sigma}$ is positive we have the thesis.
Since $\tau_{*}\rho_{\sigma}$ is differentiable and the convergence of the $F^{\prime}_{N}$ is uniform, we have that
$(N_{\sigma}f)^{\prime}(x)=\frac{d}{dx}\int_{-1}^{1}\tau_{*}(\rho_{\sigma})(x-y)f(y)dy=\int_{-1}^{1}\frac{\partial\tau_{*}(\rho_{\sigma})}{\partial x}(x-y)f(y)dy.$
Note that for all $N$ we have $F^{\prime}_{N}(x)=(\sum_{j=-N}^{N}\rho_{\sigma}(x+2j))^{\prime}=\sum_{j=-N}^{N}\rho^{\prime}_{\sigma}(x+2j)$, therefore $(\tau_{*}\rho_{\sigma})^{\prime}=\tau_{*}\rho^{\prime}_{\sigma}$, which implies the thesis.
Since $N_{\sigma}f$ is continuously differentiable, it belongs to BV([1,1])BV([-1,1]), and the variation is given by the $L^{1}$-norm of its derivative
[…]
This again follows from Young’s inequality, applied to the derivative of the convolution kernel.
Since $f\geq 0$ and $\int fdx=1$ there exists a positive Lebesgue measure set $A\subset[-1,1]$ such that $f(y)>0$ on $A$. Therefore
$N_{\sigma}f(x)=\int_{A}\sum_{j=-\infty}^{+\infty}\rho_{\sigma}(x-y+2j)f(y)dy>\min_{x\in[-1,1]}\tau_{*}\rho_{\sigma}(x)\int_{A}f(y)dy,$
since $\tau_{*}\rho_{\sigma}(x)\geq\rho_{\sigma}(x)$ for all $x\in[-1,1]$ and the minimum of $\rho_{\sigma}(x)$ on [1,1][-1,1] is $\rho_{\sigma}(1)=\frac{1}{\sigma\sqrt{2\pi}}e^{-1/2\sigma^{2}}$ we obtain the claimed lower bound. ∎
The fact that $N_{\sigma}$ is regularizing $L^{1}$ into $BV$ has important consequences
Lemma 6.3.
Let $P_{\sigma}:L^{1}([-1,1])\rightarrow L^{1}([-1,1])$, $P_{\sigma}=N_{\sigma}P$ be the annealed Perron-Frobenius operator associated to the random dynamical system with Gaussian noise of variance $\sigma$.
  • $P_{\sigma}||_{L^{1}}\leq 1$,
  • $P_{\sigma}$ maps $L^{1}([-1,1])$ into BV([1,1])BV([-1,1]),
  • $P_{\sigma}$ is a compact operator from $L^{1}([-1,1])$ to $L^{1}([-1,1])$,
  • $P_{\sigma}$ has at least a fixed point, i.e., there exists at least a stationary density.
Proof.
By Lemma 6.2 we have that
[…]
since $P||_{L^{1}}\leq 1$, as it is the Perron-Frobenius operator of a non-singular transformation.
Moreover,
[…]
therefore $P_{\sigma}$ maps a bounded sequence in $L^{1}$ into a bounded sequence in $BV$, which proves that $P_{\sigma}$ maps bounded sets in $L^{1}$ into bounded subsets of $BV$. Since the inclusion $BV([-1,1])\hookrightarrow L^{1}([-1,1])$ is compact, this implies that $P_{\sigma}:L^{1}\to L^{1}$ is compact.
From this same compactness inclusion property, we have then (following the classical Krylov-Bogolyubov argument) that the sequence
$f_{N}=\frac{1}{N}\sum_{j=0}^{N-1}P_{\sigma}^{j}\chi_{[-1,1]}$
has a subsequence that converges in $L^{1}$ to a fixed point of $P_{\sigma}$, i.e., a stationary density. ∎
Remark 6.4.
The Krylov-Bogolyubov argument ensures the existence of a fixed point, but does not guarantee uniqueness.
Definition 6.5.
Let
$\mathcal{U}_{0}=\{f\in L^{2}([-1,1])|\int fdm=0\}.$
We call $\mathcal{U}_{0}$ the vector subspace of average 0 measures. We say $P_{\sigma}$ contracts the space of average 0 functions in $L^{2}$ if
$\|P_{\sigma}^{n}|_{\mathcal{U}_{0}}\|_{L^{2}\rightarrow L^{2}}\leq C\theta^{n}$
for constants $C>0$, $0<\theta<1$.
We prove now a version of Doeblin’s argument adapted to our discussion with densities and operators.
Proposition 6.6.
Suppose that there exists $c>0$ such that for every density $f$ we have $P_{\sigma}f>c$. Then
$\|P_{\sigma}^{n}|_{\mathcal{U}_{0}}\|_{L^{1}\rightarrow L^{1}}\leq(1-2c)^{n}.$
Proof.
From Lemma 5.9, we know that $P_{\sigma}$ acts on densities. Let $h\in\mathcal{U}_{0}$ with $\|h\|_{L^{1}}=1$. Then $h$ can be written as h=fgh=f-g, where $f,g\geq 0$, $\int f\,dm=\int g\,dm$, and $\|f\|_{L^{1}}+\|g\|_{L^{1}}=1$. Hence, $\|f\|_{L^{1}}=\|g\|_{L^{1}}=\frac{1}{2}$.
By hypothesis, since $f$ and $g$ are densities with mass $\frac{1}{2}$, we have
$P_{\sigma}f\geq\frac{c}{2},\quad P_{\sigma}g\geq\frac{c}{2}\quad\text{almost everywhere}.$
Since $P_{\sigma}$ is a positive operator, we have
$P_{\sigma}h=P_{\sigma}f-P_{\sigma}g,$
and thus
$\|P_{\sigma}h\|_{L^{1}}$ $\leq\left\|P_{\sigma}f-\frac{c}{2}\right\|_{L^{1}}+\left\|P_{\sigma}g-\frac{c}{2}\right\|_{L^{1}}$
$=\int_{[-1,1]}\left|P_{\sigma}f-\frac{c}{2}\right|dm+\int_{[-1,1]}\left|P_{\sigma}g-\frac{c}{2}\right|dm.$
Since $P_{\sigma}f\geq\frac{c}{2}$ and $P_{\sigma}g\geq\frac{c}{2}$, we may drop the absolute values:
$\|P_{\sigma}h\|_{L^{1}}$ $=\int_{[-1,1]}\left(P_{\sigma}f-\frac{c}{2}\right)dm+\int_{[-1,1]}\left(P_{\sigma}g-\frac{c}{2}\right)dm$
$=\|P_{\sigma}f\|_{L^{1}}-c+\|P_{\sigma}g\|_{L^{1}}-c$
$\leq\frac{1}{2}-c+\frac{1}{2}-c=1-2c.$
This shows that
$\|P_{\sigma}|_{\mathcal{U}_{0}}\|_{L^{1}\to L^{1}}\leq 1-2c,$
and by iteration,
$\|P_{\sigma}^{n}|_{\mathcal{U}_{0}}\|_{L^{1}\to L^{1}}\leq(1-2c)^{n},$
as claimed. ∎
Proposition 6.7.
Let $\rho_{\sigma}$ be the Gaussian noise kernel. Then
$\|P_{\sigma}^{n}|_{\mathcal{U}_{0}}\|_{L^{2}\to L^{2}}\leq C\theta^{n},$
where $C=\left(\sqrt{2}+3\sqrt{2}\rho_{\sigma}(0)\right)^{1/2}$ and $\theta=(1-2c)^{1/2}\in(0,1)$.
Proof.
From the positivity of the Gaussian kernel and the definition of $P_{\sigma}$, from Lemma 6.2 we have:
$P_{\sigma}f(x)=\int_{-1}^{1}\sum_{k\in\mathbb{Z}}\rho_{\sigma}(x+2k-y)Pf(y)\,dy\geq c:=\frac{1}{\sigma\sqrt{2\pi}}e^{-1/(2\sigma^{2})},$
for all densities $f$. Thus, Proposition 6.6 applies and yields:
$\|P_{\sigma}^{n}h\|_{L^{1}}\leq(1-2c)^{n}\|h\|_{L^{1}},\quad\forall h\in\mathcal{U}_{0}.$
Now estimate the $L^{2}$ norm using interpolation:
$\|P_{\sigma}^{n}h\|_{L^{2}}^{2}$ $\leq\|P_{\sigma}^{n}h\|_{L^{1}}\cdot\|P_{\sigma}^{n}h\|_{L^{\infty}}$
$=\|P_{\sigma}^{n}h\|_{L^{1}}\cdot\|P_{\sigma}(P_{\sigma}^{n-1}h)\|_{L^{\infty}}$
$\leq\|P_{\sigma}^{n}h\|_{L^{1}}\cdot\|\tau_{*}\rho_{\sigma}*(P(P_{\sigma}^{n-1}h))^{\wedge}\|_{L^{\infty}}$
$\leq\|P_{\sigma}^{n}h\|_{L^{1}}\cdot\|\tau_{*}\rho_{\sigma}\|_{L^{\infty}}\cdot\|P(P_{\sigma}^{n-1}h)\|_{L^{1}}$
(7) $\leq(1-2c)^{n}\cdot(1+3\rho_{\sigma}(0))\cdot\|h\|_{L^{1}},$
where the last inequality uses Lemma 6.2 and that $\|P\|_{L^{1}\to L^{1}}\leq 1$.
Now normalize and pass to the operator norm:
$\|P_{\sigma}^{n}|_{\mathcal{U}_{0}}\|_{L^{2}\to L^{2}}=\sup_{\begin{subarray}{c}h\in\mathcal{U}_{0}\\ \|h\|_{L^{2}}\leq 1\end{subarray}}\|P_{\sigma}^{n}h\|_{L^{2}}\leq\sup_{\|h\|_{L^{2}}\leq 1}\left((1-2c)^{n}(1+3\rho_{\sigma}(0))\|h\|_{L^{1}}\right)^{1/2}.$
Using $\|h\|_{L^{1}}\leq\sqrt{2}\|h\|_{L^{2}}$ for $h\in\mathcal{U}_{0}\subset L^{2}$, we obtain:
$\|P_{\sigma}^{n}|_{\mathcal{U}_{0}}\|_{L^{2}\to L^{2}}\leq(1-2c)^{n/2}\left(\sqrt{2}+3\sqrt{2}\rho_{\sigma}(0)\right)^{1/2},$
as claimed. Since $c>0$, it follows that $0<\theta=\sqrt{1-2c}<1$. ∎
In Proposition 6.7, we demonstrated that $P_{\sigma}$ contracts the space of average 0 functions in $L^{2}$. We can then prove that $P_{\sigma}$ has a unique fixed point, as shown in the following proposition.
Proposition 6.8.
If $P_{\sigma}$ contracts the space of average 0 functions in $L^{2}$, then $P_{\sigma}$ has a unique fixed point.
Proof.
We prove by contradiction. Let $\mu$ and $\nu$ be stationary measures. Since $L_{\sigma}\mu=\mu$ and $L_{\sigma}\nu=\nu$ we have that $\mu$ and $\nu$ are absolutely continuous with respect to Lebesgue measure (Lemma 5.4), with densities $f$ and $g$ respectively. Now, $P_{\sigma}f=f$ and $P_{\sigma}g=g$, and since $P_{\sigma}$ contracts the space of average 0 measures, we have that for any $n$
$\|f-g\|_{L^{2}}$ $=\|P^{n}_{\sigma}f-P^{n}_{\sigma}g\|_{L^{2}}$
$=\|P^{n}_{\sigma}(f-g)\|_{L^{2}}$
$\leq C\theta^{n}\|f-g\|_{L^{2}}.$
Take $N$ such that $C\theta^{N}<1$. The inequality above implies that $\|f-g\|_{L^{1}}=0$, which in turn implies that $\mu=\nu$. Therefore, $L_{\sigma}$ has a unique stationary measure, it follows that $P_{\sigma}$ has a unique fixed point. ∎
7. Fourier theory and Galerkin approximation
In this section, we formalize the use of Fourier analysis to study the action of the smoothing operator $N_{\sigma}$. We define appropriate Banach spaces of sequences (e.g., $\ell^{\infty}$, $\ell^{2}$, and exponentially decaying sequences $\ell^{\exp}_{\sigma}$), and show how the Fourier transform maps functions on [1,1][-1,1] into these spaces.
We prove that $N_{\sigma}$ is diagonalized by the Fourier basis, with eigenvalues given by a diagonal operator $D_{\sigma}$ (Proposition 7.6), and that the entries of $D_{\sigma}$ decay exponentially fast (Theorem 7.8). This result justifies the use of Galerkin-type finite-dimensional approximations, since high-frequency modes are exponentially suppressed.
Definition 7.1.
We define the Fourier transform on [1,1][-1,1] as the map
$\mathcal{F}(f)[k]:=\frac{1}{2}\int_{-1}^{1}f(x)e^{-k\pi ix}\,dx,$
for all $k\in\mathbb{Z}$. The number $\mathcal{F}(f)[k]$ is called the $k$-th Fourier coefficient of $f$.
Definition 7.2.
Let $\ell^{\infty}$ denote the space of bounded bilateral sequences:
$\ell^{\infty}=\{(a_{k})_{k\in\mathbb{Z}}:|a_{k}|\leq C\text{ for some }C>0\}.$
Definition 7.3.
For $\sigma>0$, define the space of exponentially decaying sequences:
$\ell^{\exp}_{\sigma}=\left\{(a_{k})_{k\in\mathbb{Z}}:|a_{k}|\leq Ce^{-\frac{\sigma^{2}|k|}{2}}\text{ for some }C>0\right\}.$
Definition 7.4.
Define $\ell^{2}$, the space of square-summable sequences:
$\ell^{2}=\left\{(a_{k})_{k\in\mathbb{Z}}:\sum_{k\in\mathbb{Z}}|a_{k}|^{2}<\infty\right\},$
endowed with the norm $\|(a_{k})\|_{\ell^{2}}:=\left(\sum|a_{k}|^{2}\right)^{1/2}$. It forms a Hilbert space with respect to the inner product $(a_{k}),(b_{k})>=\sum a_{k}\bar{b}_{k}$.
Lemma 7.5.
The Fourier transform $\mathcal{F}$ on [1,1][-1,1] satisfies the following properties:
  1. (1)
    $\mathcal{F}:L^{1}([-1,1])\to\ell^{\infty}$ is continuous with operator norm $\|\mathcal{F}\|=1$.
  2. (2)
    $\mathcal{F}$ is invertible on $\ell^{\exp}_{\sigma}$,
  3. (3)
    $\mathcal{F}:L^{2}([-1,1])\to\ell^{2}$ is an isometry, and the family $\{e^{-k\pi ix}\}_{k\in\mathbb{Z}}$ forms an orthonormal basis of $L^{2}([-1,1])$.
Proof.
We only give a short proof of item 2. Since $\{a_{k}\}_{k\in\mathbb{Z}}$ is in $\ell^{\exp}_{\sigma}$, we have
$\left|\displaystyle\sum_{k=-\infty}^{+\infty}a_{k}e^{\pi ikx}\right|$ $\leq\displaystyle\sum_{k=-\infty}^{+\infty}\left|a_{k}\right|<\infty,$
and from Weierstrass’s M-test Theorem
$\mathcal{F}^{-1}(\{a_{k}\})=\displaystyle\sum_{k=-\infty}^{+\infty}a_{k}e^{\pi ikx}$
converges uniformly to an analytic function. ∎
The operator $N_{\sigma}$ has an explicit representation that commutes with the Fourier transform.
Proposition 7.6.
There exists an operator $D_{\sigma}:\ell_{\infty}\to\ell_{\infty}$ such that
$D_{\sigma}\mathcal{F}=\mathcal{F}N_{\sigma}.$
Proof.
We compute the $k$-th Fourier coefficient of $N_{\sigma}f$:
$\mathcal{F}(N_{\sigma}f)[k]$ $=\frac{1}{2}\int_{-1}^{1}\left(\sum_{j\in\mathbb{Z}}\int_{-1}^{1}\rho_{\sigma}(x+2j-y)f(y)\,dy\right)e^{-k\pi ix}\,dx$
$=\frac{1}{2}\sum_{j\in\mathbb{Z}}\int_{-1}^{1}\int_{-1}^{1}\rho_{\sigma}(x+2j-y)f(y)e^{-k\pi ix}\,dydx$
$=\frac{1}{2}\sum_{j\in\mathbb{Z}}\int_{-1}^{1}\int_{-1}^{1}\rho_{\sigma}(z+2j)f(y)e^{-k\pi i(z+y)}\,dzdy\quad(z=x-y)$
$=\frac{1}{2}\left(\sum_{j\in\mathbb{Z}}\int_{-1}^{1}\rho_{\sigma}(z+2j)e^{-k\pi iz}\,dz\right)\cdot\mathcal{F}(f)[k].$
Define the diagonal operator $D_{\sigma}$ on $\ell_{\infty}$ by
$D_{\sigma}[k,k]:=\frac{1}{2}\int_{-\infty}^{+\infty}\rho_{\sigma}(z)e^{-k\pi iz}\,dz,$
so that $\mathcal{F}(N_{\sigma}f)[k]=D_{\sigma}[k,k]\cdot\mathcal{F}(f)[k]$. Hence,
$\mathcal{F}(N_{\sigma}f)=D_{\sigma}\mathcal{F}(f),$
as claimed. ∎
We observe now that the coefficients $D[k,k]$ can be obtained by a classical computation, i.e. that
$\int_{-\infty}^{+\infty}\rho_{\sigma}(u)e^{-k\pi iu}du=e^{-\frac{\sigma^{2}(k\pi)^{2}}{2}}.$
This is the Fourier transform on $\mathbb{R}$ of a Gaussian [36, Theorem 1.4, Chapter 5]
Corollary 7.7.
Let $D_{\sigma}$ from Proposition 7.6 then
$D[k,k]=\int_{-\infty}^{+\infty}\rho_{\sigma}(z)e^{-k\pi iz}dz=e^{-\frac{\sigma^{2}k^{2}\pi^{2}}{2}}.$
In particular, the entries $D[k,k]$ converge to 0 exponentially fast as $|k|$ goes to infinity.
Theorem 7.8.
Let $\sigma>0$, and let $D_{\sigma}:\ell_{\infty}\to\ell_{\infty}$ be the diagonal operator defined by
$D_{\sigma}[k,k]=\int_{-\infty}^{+\infty}\rho_{\sigma}(z)e^{-k\pi iz}\,dz=e^{-\frac{\sigma^{2}k^{2}\pi^{2}}{2}}.$
Then:
  1. (1)
    The image of $D_{\sigma}$ lies in the space of exponentially decaying sequences
    $\ell^{\exp}_{\sigma}:=\left\{(a_{k})_{k\in\mathbb{Z}}\;\middle|\;\exists C>0\text{ such that }|a_{k}|\leq Ce^{-\frac{\sigma^{2}|k|}{2}}\text{ for all }k\in\mathbb{Z}\right\}.$
  2. (2)
    In particular, for any $f\in L^{1}([-1,1])$, the Fourier coefficients of $N_{\sigma}f$ satisfy
    $\mathcal{F}(N_{\sigma}f)=D_{\sigma}\mathcal{F}(f)\in\ell^{\exp}_{\sigma},$
    i.e., $N_{\sigma}f$ has exponentially decaying Fourier coefficients.
Proof.
Let $(b_{k})_{k\in\mathbb{Z}}\in\ell_{\infty}$, so that $|b_{k}|\leq C$ for all $k$. Then
$D_{\sigma}(b_{k})=(e^{-\frac{\sigma^{2}\pi^{2}k^{2}}{2}}b_{k})_{k\in\mathbb{Z}}.$
Since $e^{-\frac{\sigma^{2}\pi^{2}k^{2}}{2}}\leq e^{-\frac{\sigma^{2}|k|}{2}}$, we have
$|D_{\sigma}(b_{k})|\leq Ce^{-\frac{\sigma^{2}|k|}{2}},$
so $D_{\sigma}(b_{k})\in\ell^{\exp}_{\sigma}$, as claimed. ∎
Corollary 7.9.
The stationary density $f_{\sigma}$ is unique and real-analytic.
Proof.
By Proposition 6.8 the annealed Perron–Frobenius operator $P_{\sigma}$ has a unique fixed point $f_{\sigma}$ in $L^{1}([{-1},1])$, corresponding to the unique stationary density of the random dynamical system.
Moreover, since $f_{\sigma}=P_{\sigma}f_{\sigma}=N_{\sigma}Pf_{\sigma}$, and $Pf_{\sigma}\in L^{1}$, Theorem 7.8 implies that $f_{\sigma}$ has exponentially decaying Fourier coefficients. Hence, $f_{\sigma}\in\ell^{\exp}_{\sigma}$, and its Fourier series converges uniformly to a real-analytic function.
Therefore, $f_{\sigma}$ is real-analytic on [1,1][-1,1]. ∎
8. Fourier Approximation
In this section, we show how to approximate the fixed point of the annealed Perron-Frobenius operator $P_{\sigma}$ using the Fourier basis.
Since $P_{\sigma}$ acts on functions, our goal is to approximate its fixed point—a function—via a finite-dimensional computation. To do this, we project $P_{\sigma}$ onto a finite-dimensional subspace spanned by the first $2k+1$ Fourier modes. This allows us to work with a truncated version of the operator and compute a numerical approximation of the fixed point in coefficient space. However, note that recovering the full analytic function from its truncated Fourier representation is not always possible.
To implement this approximation, we define a projection operator $\pi_{k}$ that truncates the Fourier expansion to the first $2k+1$ terms.
Definition 8.1.
Let $\pi_{k}:L^{1}([-1,1])\to L^{2}([-1,1])$ be defined by
$\pi_{k}f(x)=\sum_{j=-k}^{k}\left(\frac{1}{2}\int_{-1}^{1}f(x)e^{-j\pi ix}dx\right)e^{j\pi ix}=\sum_{j=-k}^{k}\mathcal{F}(f)[j]e^{j\pi ix}.$
This is the Galerkin projection onto the Fourier modes of order at most $k$.
To handle truncation directly in Fourier coefficient space, we define the projection $\widetilde{\pi}_{k}$ on sequences:
Definition 8.2.
Let $\widetilde{\pi}_{k}:\ell_{\infty}\to\ell_{\infty}$ be defined by
$\widetilde{\pi}_{k}((a_{j})_{j\in\mathbb{Z}})=(b_{j})_{j\in\mathbb{Z}},\quad\text{where }b_{j}=\begin{cases}a_{j},&|j|\leq k,\\ 0,&|j|>k.\end{cases}$
The following lemma summarizes key algebraic properties of $\pi_{k}$, $\widetilde{\pi}_{k}$, and their relation to the diagonal operator $D_{\sigma}$ defined earlier.
Lemma 8.3.
Let $\pi_{k}$, $\widetilde{\pi}_{k}$, and $D_{\sigma}$ be as above. Then:
  1. (1)
    $\pi_{k}^{n}=\pi_{k}$ and $\widetilde{\pi}_{k}^{n}=\widetilde{\pi}_{k}$ for all $n\in\mathbb{N}$ (they are projections).
  2. (2)
    If $k<k^{\prime}$, then $\pi_{k}\pi_{k^{\prime}}=\pi_{k}$.
  3. (3)
    $D_{\sigma}$ commutes with both $\widetilde{\pi}_{k}$ and $1-\widetilde{\pi}_{k}$.
  4. (4)
    The following identity holds:
    $\mathcal{F}^{-1}\widetilde{\pi}_{k}\mathcal{F}=\pi_{k}.$
Recall that $D_{\sigma}\mathcal{F}=\mathcal{F}N_{\sigma}$ and that the diagonal entries of $D_{\sigma}$ decay exponentially. We now quantify how close $N_{\sigma}$ is to its truncated version.
Theorem 8.4 (Tail Estimate for $N_{\sigma}$).
Let $\Gamma_{\sigma,k}:=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{\sigma^{2}k^{2}\pi^{2}}{2}}$. Then:
$\|N_{\sigma}(1-\pi_{k})\|_{L^{1}\to L^{2}}\leq\Gamma_{\sigma,k},\quad\|(1-\pi_{k})N_{\sigma}\|_{L^{1}\to L^{2}}\leq\Gamma_{\sigma,k}.$
Sketch of Proof.
Let $f\in L^{1}([-1,1])$. By Parseval and the exponential decay of the diagonal entries $D[k,k]$ (see Lemma 7.7), we have
$\|N_{\sigma}(1-\pi_{k})f\|_{L^{2}}^{2}=\sum_{|j|>k}|D[j,j]|^{2}|\mathcal{F}(f)[j]|^{2}\leq\Gamma_{\sigma,k}^{2}\|f\|_{L^{1}}^{2}.$
Next, we show that the projected discretization of $P_{\sigma}$ coincides with that of the individual components:
Lemma 8.5.
The following identity holds:
$\pi_{k}N_{\sigma}\pi_{k}P\pi_{k}=\pi_{k}P_{\sigma}\pi_{k}.$
Proof.
Follows directly from the identities $P_{\sigma}=N_{\sigma}P$ and $\pi_{k}^{2}=\pi_{k}$. ∎
While $N_{\sigma}$ is a contraction on $L^{1}$, the discretized version $\pi_{k}N_{\sigma}$ may, in principle, increase the norm. We now bound this growth:
Lemma 8.6 (Norm Bound for Projected Operator).
The projected operator satisfies
$\|\pi_{k}N_{\sigma}\|_{L^{1}\to L^{1}}\leq 1+\Gamma_{\sigma,k}.$
Proof.
We use the identity $\pi_{k}N_{\sigma}=N_{\sigma}\pi_{k}$, which follows from the commutation relations between $D_{\sigma}$, $\widetilde{\pi}_{k}$, and $\mathcal{F}$ (see Lemma 8.3). Then
$\|\pi_{k}N_{\sigma}f\|_{L^{1}}$ $\leq\|N_{\sigma}f\|_{L^{1}}+\|(1-\pi_{k})N_{\sigma}f\|_{L^{1}}$
$\leq\|f\|_{L^{1}}+\|(1-\pi_{k})N_{\sigma}\|_{L^{1}\to L^{2}}\|f\|_{L^{1}}$
$\leq(1+\Gamma_{\sigma,k})\|f\|_{L^{1}}.$
Remark 8.7.
Together, these results justify the use of the finite-rank approximation $\pi_{k}P_{\sigma}\pi_{k}$ as a computable and controlled approximation of $P_{\sigma}$, with a well-quantified truncation error.
Next, we compare the $L^{1}$ norm of the fixed point of $P_{\sigma}$ with the norm of its Galerkin projection.
Lemma 8.8.
Let $f_{\sigma}$ be the fixed point of $P_{\sigma}$. Then
$\|\pi_{k}f_{\sigma}\|_{L^{1}}=\|\pi_{k}N_{\sigma}Pf_{\sigma}\|_{L^{1}}\leq(1+\Gamma_{\sigma,k})\|f_{\sigma}\|_{L^{1}},$
where $\Gamma_{\sigma,k}:=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{\sigma^{2}k^{2}\pi^{2}}{2}}$.
We now define the discretized version of $P_{\sigma}$, which we will use in our approximation scheme.
Definition 8.9.
The discretized annealed Perron-Frobenius operator is defined as
$P_{\sigma,k}:=\pi_{k}N_{\sigma}\pi_{k}P\pi_{k}=\pi_{k}N_{\sigma}P\pi_{k}=\pi_{k}P_{\sigma}\pi_{k}:L^{1}\to L^{2}.$
Note that if $f\in\mathcal{U}_{0}$, the space of zero-average functions in $L^{2}([-1,1])$, then
$\mathcal{F}(f)[0]=\frac{1}{2}\int_{-1}^{1}f(x)dx=0.$
We now state our main theorem, a variation of Theorem 3.4 from [20], adapted to our context. It provides a rigorous bound on the error between the true fixed point $f_{\sigma}$ and an approximate fixed point obtained from a finite-dimensional Galerkin scheme.
Theorem 8.10.
Let $f_{\sigma}$ be the unique fixed point of $P_{\sigma}$, and let $P_{\sigma,k}$ be its Galerkin approximation. Suppose there exist $n\in\mathbb{N}$ and $C_{n}>0$ such that
$\|P_{\sigma,k}^{n}|_{\mathcal{U}_{0}}\|_{L^{2}\to L^{2}}\leq C_{n}<1,$
and that for $0\leq i<n$, we have bounds $1\leq\|P_{\sigma,k}^{i}|_{\mathcal{U}_{0}}\|_{L^{2}\to L^{2}}\leq C_{i}$. Let $g$ be a trigonometric polynomial of degree at most $k$ such that $\|P_{\sigma,k}g-g\|_{L^{2}}\leq\epsilon$. Then
$\|f_{\sigma}-g\|_{L^{2}}\leq\frac{1}{1-C_{n}}\sum_{i=0}^{n-1}C_{i}\left((1+\Gamma_{\sigma,k}+\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})})\Gamma_{\sigma,k}\|f_{\sigma}\|_{L^{1}}+\epsilon\right).$
Remark 8.11.
All constants involved in the theorem can be rigorously computed or estimated with the aid of a computer. In particular, the contraction bound $C_{n}$, the intermediate norms $C_{i}$, and the residual $\|P_{\sigma,k}g-g\|_{L^{2}}$ are all computable. The choice of the approximation $g$ is crucial: we typically use a numerically computed fixed point of $P_{\sigma,k}$, post-processed to enforce symmetry and improve approximation of the zero-frequency component.
To prove Theorem 8.10, we first recall two lemmas adapted from [20], which control the behavior of powers of the discretized operator $P_{\sigma,k}$ restricted to the subspace $\mathcal{U}_{0}$.
Lemma 8.12 ([20]).
Let $C_{i}:=\|P_{\sigma,k}^{i}|_{\mathcal{U}_{0}}\|_{L^{2}\rightarrow L^{2}}$ for $i\in\mathbb{N}_{0}$, and suppose that $C_{n}<1$ for some $n\in\mathbb{N}$. Then:
  1. (1)
    $\sum_{i=0}^{\infty}C_{i}\leq\frac{1}{1-C_{n}}\sum_{i=0}^{n-1}C_{i}$,
  2. (2)
    There exist constants $C>0$ and $\theta\in(0,1)$ such that $\|P_{\sigma,k}^{i}|_{\mathcal{U}_{0}}\|_{L^{2}\to L^{2}}\leq C\theta^{i}$ for all $i$.
Proof.
(1) Write $i=qn+r$ with $0\leq r<n$. Using the invariance $P_{\sigma,k}(\mathcal{U}_{0})\subset\mathcal{U}_{0}$, we get:
$\|P_{\sigma,k}^{i}|_{\mathcal{U}_{0}}\|\leq\|P_{\sigma,k}^{n}\|^{q}\cdot\|P_{\sigma,k}^{r}\|\leq C_{n}^{q}C_{r}.$
Summing over $i$, we reorganize by blocks of length $n$:
$\sum_{i=0}^{\infty}C_{i}\leq\sum_{q=0}^{\infty}C_{n}^{q}\sum_{r=0}^{n-1}C_{r}=\left(\sum_{q=0}^{\infty}C_{n}^{q}\right)\sum_{r=0}^{n-1}C_{r}=\frac{1}{1-C_{n}}\sum_{i=0}^{n-1}C_{i}.$
(2) From the previous bound,
$\|P_{\sigma,k}^{i}\|\leq C_{n}^{\lfloor i/n\rfloor}\cdot\max_{r<n}C_{r}\leq C_{n}^{i/n-1}\cdot\max_{r<n}C_{r},$
so the desired exponential decay holds with $\theta=C_{n}^{1/n}$ and $C=\max_{r<n}C_{r}/C_{n}$. ∎
We now relate the size of a vector in $\mathcal{U}_{0}$ to the residual of its image under $P_{\sigma,k}$.
Lemma 8.13.
Let $C_{i}:=\|P_{\sigma,k}^{i}|_{\mathcal{U}_{0}}\|_{L^{2}\to L^{2}}$ with $C_{n}<1$ for some $n\in\mathbb{N}$. Then for any $v\in\mathcal{U}_{0}$,
$\|v\|_{L^{2}}\leq\frac{1}{1-C_{n}}\sum_{i=0}^{n-1}C_{i}\|P_{\sigma,k}v-v\|_{L^{2}}.$
Proof.
We telescope:
$v=(P_{\sigma,k}v-v)+(P_{\sigma,k}^{2}v-P_{\sigma,k}v)+\cdots+(P_{\sigma,k}^{n}v-P_{\sigma,k}^{n-1}v)-P_{\sigma,k}^{n}v.$
Taking $L^{2}$ norms:
$\|v\|_{L^{2}}\leq\sum_{i=0}^{n-1}\|P_{\sigma,k}^{i}(P_{\sigma,k}v-v)\|_{L^{2}}+\|P_{\sigma,k}^{n}v\|_{L^{2}}.$
By Lemma 8.12(2), $\|P_{\sigma,k}^{n}v\|\to 0$ as $n\to\infty$. So:
$\|v\|_{L^{2}}\leq\sum_{i=0}^{\infty}C_{i}\|P_{\sigma,k}v-v\|_{L^{2}}\leq\frac{1}{1-C_{n}}\sum_{i=0}^{n-1}C_{i}\|P_{\sigma,k}v-v\|_{L^{2}},$
as claimed. ∎
Lemma 8.14.
Let $\rho_{\sigma}(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{x^{2}}{2\sigma^{2}}}$ be the Gaussian density, and let $\tau_{*}\rho_{\sigma}(x):=\sum_{j\in\mathbb{Z}}\rho_{\sigma}(x+2j)$ be its $2$-periodic summation. Then
$\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}=\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}=\frac{1}{2\sigma\sqrt{\pi}}.$
Proof.
We prove both directions of the identity.
(1) Lower bound: By orthogonality of the supports (modulo period shifts),
$\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}^{2}=\int_{-1}^{1}\left|\sum_{j\in\mathbb{Z}}\rho_{\sigma}(x+2j)\right|^{2}dx\geq\sum_{j\in\mathbb{Z}}\int_{-1}^{1}|\rho_{\sigma}(x+2j)|^{2}dx=\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}^{2}.$
(2) Upper bound: Parseval’s identity for Fourier series on [1,1][-1,1] gives:
$\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}^{2}=2\sum_{k\in\mathbb{Z}}|\mathcal{F}(\tau_{*}\rho_{\sigma})[k]|^{2}.$
The Fourier coefficients of $\tau_{*}\rho_{\sigma}$ are computed by extending the integral to $\mathbb{R}$:
$\mathcal{F}(\tau_{*}\rho_{\sigma})[k]=\frac{1}{2}\int_{\mathbb{R}}\rho_{\sigma}(x)e^{-k\pi ix}dx=\frac{1}{2}e^{-\frac{\sigma^{2}\pi^{2}k^{2}}{2}}.$
Hence,
$\|\tau_{*}\rho_{\sigma}\|_{L^{2}([-1,1])}^{2}=\sum_{k\in\mathbb{Z}}e^{-\sigma^{2}\pi^{2}k^{2}}=\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}^{2}.$
(3) Classical integral identity: It is well-known [34] that
$\int_{-\infty}^{+\infty}e^{-ax^{2}}dx=\sqrt{\frac{\pi}{a}}\quad\text{for }a>0.$
Applying this with $a=1/\sigma^{2}$,
$\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}^{2}=\int_{-\infty}^{\infty}\left(\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{x^{2}}{2\sigma^{2}}}\right)^{2}dx=\frac{1}{\sigma^{2}(2\pi)}\int_{-\infty}^{\infty}e^{-\frac{x^{2}}{\sigma^{2}}}dx=\frac{1}{2\sigma\sqrt{\pi}}.$
This completes the proof. ∎
Proof of Theorem 8.10.
Let $v=f_{\sigma}-g$. Then
[…]
Using linearity and triangle inequality:
$\|P_{\sigma,k}v-v\|_{L^{2}}\leq\|P_{\sigma,k}f_{\sigma}-f_{\sigma}\|_{L^{2}}+\|P_{\sigma,k}g-g\|_{L^{2}}=\delta+\epsilon,$
where $\epsilon=\|P_{\sigma,k}g-g\|_{L^{2}}$. It remains to bound $\delta=\|P_{\sigma,k}f_{\sigma}-f_{\sigma}\|_{L^{2}}$.
Using Lemma 8.5, we write
$P_{\sigma,k}f_{\sigma}-f_{\sigma}=(\pi_{k}N_{\sigma}\pi_{k}P\pi_{k}-N_{\sigma}P)f_{\sigma}=(1-\pi_{k})N_{\sigma}P\pi_{k}f_{\sigma}+N_{\sigma}P(1-\pi_{k})f_{\sigma}.$
For the first term, by Theorem 8.4 and Lemma 8.8:
$\|(1-\pi_{k})N_{\sigma}P\pi_{k}f_{\sigma}\|_{L^{2}}\leq\Gamma_{\sigma,k}\|\pi_{k}f_{\sigma}\|_{L^{1}}\leq\Gamma_{\sigma,k}(1+\Gamma_{\sigma,k})\|f_{\sigma}\|_{L^{1}}.$
For the second term, using Lemma 8.14 and Young’s inequality:
$\|N_{\sigma}P(1-\pi_{k})f_{\sigma}\|_{L^{2}}\leq\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}\cdot\Gamma_{\sigma,k}\|f_{\sigma}\|_{L^{1}}.$
Combining both:
$\delta=\|P_{\sigma,k}f_{\sigma}-f_{\sigma}\|_{L^{2}}\leq(1+\Gamma_{\sigma,k}+\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})})\Gamma_{\sigma,k}\|f_{\sigma}\|_{L^{1}}.$
Plugging back:
$\|f_{\sigma}-g\|_{L^{2}}\leq\frac{1}{1-C_{n}}\sum_{i=0}^{n-1}C_{i}\left[(1+\Gamma_{\sigma,k}+\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})})\Gamma_{\sigma,k}\|f_{\sigma}\|_{L^{1}}+\epsilon\right].\qed$
9. Proof of Main Result 1
This section establishes the abstract computability of the stationary density for the annealed system under Gaussian noise, and proves Main Result 1. The underlying idea is that the a priori estimate on the mixing rates in Proposition 6.7 allows us to bound uniformly the mixing rates of the operators $P_{\sigma,k}$ for all $k>K$, which in turn allows us to prove that the approximation error on the stationary measure goes to 0 as $k$ goes to infinity.
Main result 1.
Let $T$ be a non-singular map of the interval, and consider the random dynamical system:
(8) $X_{n+1}=\tau(T(X_{n})+\Omega_{\sigma}(n)),$
where $\Omega_{\sigma}(n)$ are i.i.d. Gaussian random variables with standard deviation $\sigma$, and $\tau$ is the periodic boundary condition (Definition 3.2).
Then, the stationary density can be approximated in $L^{2}$, and Birkhoff averages of any observable $\phi\in L^{2}$ can be enclosed to arbitrary precision.
This computability result relies crucially on the strong regularization provided by the Gaussian noise (see Proposition 6.7). Our strategy proceeds by controlling the discretization error of the annealed transfer operator $P_{\sigma}$, and showing convergence of the approximated fixed point $f_{\sigma,K}$ to the true fixed point $f_{\sigma}$ in $L^{2}$ norm.
9.1. Norm Bounds for the Discretized Operator
We first bound the $L^{1}$ norm of iterates of the discretized operator.
Lemma 9.1.
Let $P_{\sigma,K}$ be the discretized annealed operator. Then for all $i\in\mathbb{N}$:
$\|P^{i}_{\sigma,K}\|_{L^{1}\to L^{1}}\leq(1+\Gamma_{\sigma,K})^{i}(2K+1).$
Proof.
Using $\pi_{K}^{2}=\pi_{K}$ (Lemma 8.3), we write:
$P^{i}_{\sigma,K}=(\pi_{K}N_{\sigma}P)^{i}\pi_{K}.$
By Lemma 8.6, $\|\pi_{K}N_{\sigma}\|_{L^{1}\to L^{1}}\leq 1+\Gamma_{\sigma,K}$, and since $\|P\|_{L^{1}\to L^{1}}=1$, the composition has norm at most $1+\Gamma_{\sigma,K}$. Lastly, since $|\mathcal{F}(g)[j]|\leq\|g\|_{L^{1}}$, we estimate:
$\|\pi_{K}g\|_{L^{1}}\leq\sum_{j=-K}^{K}|\mathcal{F}(g)[j]|\leq(2K+1)\|g\|_{L^{1}}.$
Combining both gives the result. ∎
Lemma 9.2.
Let $P_{\sigma}$ be the annealed transfer operator. Then, there exist $K,N\in\mathbb{N}$ such that for all $k>K$:
$\|P^{N}_{\sigma,k}|_{\mathcal{U}_{0}}\|_{L^{2}}\leq 3/4.$
Sketch of proof.
Choose NN such that $\|P_{\sigma}^{N}|_{\mathcal{U}_{0}}\|_{L^{2}}\leq 1/2$ (from Proposition 6.7). Then for any $g\in\mathcal{U}_{0}$:
$\|P_{\sigma,k}^{N}g\|_{L^{2}}$ $\leq\|P_{\sigma}^{N}g\|_{L^{2}}+\|P_{\sigma}^{N}g-P_{\sigma,k}^{N}g\|_{L^{2}}$
$\leq\frac{1}{2}+\text{(discretization error)}.$
The error is expanded using a telescoping sum:
$\sum_{i=0}^{N-1}\|P_{\sigma}^{i}(P_{\sigma}-P_{\sigma,k})P_{\sigma,k}^{N-i-1}g\|_{L^{2}},$
and each term is bounded using Lemma 9.1, Proposition 8.10, and exponential decay of $\Gamma_{\sigma,k}$. The sum is ultimately controlled by:
$\sum_{i=0}^{N-1}C_{i}(1+\Gamma_{\sigma,k})^{N}(2K+1)\Gamma_{\sigma,k},$
which can be made less than $1/4$ for large enough $K$. ∎
Lemma 9.3.
For all $i\in\mathbb{N}$, we have:
$\|P_{\sigma,K}^{i}\|_{L^{2}}\leq\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}(1+\Gamma_{\sigma,K})^{i-1}(2K+1)\sqrt{2}.$
Sketch of proof.
The bound follows from combining:
  • Young’s inequality for convolution with $\tau_{*}\rho_{\sigma}$ (see Lemma 6.2),
  • the fact that $\|N_{\sigma}\|_{L^{1}\to L^{2}}=\|\rho_{\sigma}\|_{L^{2}(\mathbb{R})}$,
  • the estimate $\|\pi_{K}g\|_{L^{1}}\leq(2K+1)\|g\|_{L^{1}}$,
  • and Lemma 9.1 for controlling $L^{1}$ norms of iterates.
Theorem 9.4.
Let $f_{\sigma}$ be the fixed point of $P_{\sigma}$, and $f_{\sigma,K}$ the fixed point of $P_{\sigma,K}$. Then for any $\epsilon>0$, there exists $K$ such that:
$\|f_{\sigma}-f_{\sigma,K}\|_{L^{2}}<\epsilon.$
Proof.
Apply Theorem 8.10 with the error term
$(1+\Gamma_{\sigma,K}+\|\rho_{\sigma}\|_{L^{2}})\Gamma_{\sigma,K}\|f_{\sigma}\|_{L^{1}},$
multiplied by the norm bounds from Lemma 9.3 and the number of terms NN in Lemma 9.2. Since $\Gamma_{\sigma,K}\to 0$ exponentially as $K\to\infty$, we can ensure the full expression is less than $\epsilon$. ∎
Corollary 9.5.
Let $\phi$ be an observable in $L^{2}(m)$; the Birkhoff average
$\frac{1}{N}\sum_{n=0}^{N-1}\phi(X_{n})$
converges almost surely to $\int\phi\,f_{\sigma}dm$, Then for any $\epsilon>0$, there exists $K$ such that:
$\left|\int\phi\,f_{\sigma}dm-\int\phi\,f_{\sigma,K}dm\right|<\epsilon.$
Proof.
Due to uniqueness of the stationary measure (Proposition 6.8) the convergence of the Birkhoff average follows from Birkhoff Ergodic Theorem 5.12; by the Cauchy-Schwarz inequality we have that
$\left|\int\phi\,f_{\sigma}dm-\int\phi\,f_{\sigma,K}dm\right|\leq\|\phi\|_{L^{2}}||\cdot\|f_{\sigma}-f_{\sigma,K}\|_{L^{2}}$
and the thesis follows from Theorem 9.4. ∎
10. Main Result 2: the needed estimates
We now apply the theory developed in previous sections to compute a rigorous enclosure of the Lyapunov exponent for the noisy dynamical system
(9) $X_{n+1}=\tau(T(X_{n})+\Omega_{\sigma}(n)),$
where $\Omega_{\sigma}(n)$ are i.i.d. Gaussian random variables with standard deviation $\sigma$, and $\tau$ is the periodic boundary projection.
Our main result will be to estimate the Lyapunov exponent
$\lambda(\alpha,\beta,\sigma)=\int_{-1}^{1}\ln|T^{\prime}_{\alpha,\beta}(x)|f_{\sigma}(x)\,dx,$
where $f_{\sigma}$ is the stationary density of the annealed Perron-Frobenius operator.
10.1. Integrability of the Observable
Lemma 10.1.
Since $\ln|T^{\prime}_{\alpha,\beta}|\in L^{1}([-1,1])$, $\mu_{\sigma}=f_{\sigma}dm$ and $f_{\sigma}$ is bounded, we have
$\ln|T^{\prime}_{\alpha,\beta}|\in L^{1}(\mu_{\sigma}).$
Moreover, the constant
$\Upsilon:=\|\ln|T^{\prime}_{\alpha,\beta}|\|_{L^{2}([-1,1])}=\sqrt{2}\left((\ln(\alpha(1+\beta))-(\alpha-1))^{2}+(\alpha-1)^{2}\right)^{1/2}$
coincides with the $L^{2}$-norm of $\ln|T^{\prime}_{\alpha,\beta}|$ and will be used in later bounds.
Proof.
Using Hölder’s inequality:
$\int_{-1}^{1}\ln|T^{\prime}_{\alpha,\beta}|\,d\mu_{\sigma}=\int_{-1}^{1}\ln|T^{\prime}_{\alpha,\beta}|f_{\sigma}\,dm\leq\|\ln|T^{\prime}_{\alpha,\beta}|\|_{L^{2}}\|f_{\sigma}\|_{L^{2}}.$
Since $f_{\sigma}$ is bounded, it lies in $L^{2}$, and we compute
$\|\ln|T^{\prime}_{\alpha,\beta}|\|_{L^{2}}^{2}$ $=2\int_{0}^{1}\left(\ln(\alpha(1+\beta))+(\alpha-1)\ln x\right)^{2}dx$
$=2\left(\ln^{2}(\alpha(1+\beta))-2\ln(\alpha(1+\beta))(\alpha-1)+2(\alpha-1)^{2}\right)$
$=2\left((\ln(\alpha(1+\beta))-(\alpha-1))^{2}+(\alpha-1)^{2}\right),$
which gives the expression for $\Upsilon$ as stated. ∎
10.2. Approximating the Lyapunov Exponent
Definition 10.2.
Let $f_{\sigma,k,s}$ be a symmetrized trigonometric polynomial approximating $f_{\sigma}$, we define
$\lambda_{s}(\alpha,\beta,\sigma,k)=\int_{-1}^{1}\ln|T^{\prime}_{\alpha,\beta}(x)|f_{\sigma,k,s}(x)\,dx.$
Theorem 10.3.
Let $f_{\sigma,k,s}$ be a symmetrized trigonometric polynomial approximating $f_{\sigma}$. Then
$|\lambda(\alpha,\beta,\sigma)-\lambda_{s}(\alpha,\beta,\sigma,k)|\leq\Upsilon\|f_{\sigma}-f_{\sigma,k,s}\|_{L^{2}},$
where $\Upsilon=\sqrt{2}((\ln(\alpha(1+\beta))-(\alpha-1))^{2}+(\alpha-1)^{2})^{1/2}$.
Proof.
By Cauchy-Schwarz,
$\left|\int_{-1}^{1}\ln|T^{\prime}|(f_{\sigma}-f_{\sigma,k,s})\,dx\right|\leq\|\ln|T^{\prime}|\|_{L^{2}}\|f_{\sigma}-f_{\sigma,k,s}\|_{L^{2}}=\Upsilon\|f_{\sigma}-f_{\sigma,k,s}\|_{L^{2}}.$
10.3. Fourier Expansion of the Observable
Lemma 10.4.
The Fourier coefficients of $\ln|T^{\prime}_{\alpha,\beta}|$ on [1,1][-1,1] are
$\mathcal{F}[0]$ $=\ln(\alpha(1+\beta))-(\alpha-1),$
$\mathcal{F}[j]$ $=-\frac{(\alpha-1)}{j\pi}\int_{0}^{j\pi}\frac{\sin(x)}{x}dx,\quad j\neq 0.$
Remark 10.5.
Note that to compute the Fourier expansion of $\ln|T^{\prime}_{\alpha,\beta}|$ we only need to compute rigorously $\int_{0}^{j\pi}\frac{1}{x}\sin(x)\,dx$. The procedure we use to compute $\int_{0}^{j\pi}\frac{1}{x}\sin(x)\,dx$ combines Taylor models and power series expansions, ensuring both rigor and numerical efficiency:
  • For intervals $l\pi,(l+1)\pi]$ where $l\geq 1$, we use high-order Taylor models to approximate the integrand and compute verified bounds for the integral.
  • Near the singularity at $x=0$, we rely on the power series expansion of $\frac{\sin x}{x}$, integrating term by term to compute the integral over $0,\pi]$ with controlled truncation error.
All calculations are implemented in interval arithmetic, guaranteeing that the computed bounds are mathematically rigorous.
10.4. Computation Strategy
Let $\widetilde{f}_{\sigma,k,s}$ be the approximation of $f_{\sigma}$ on $[0,1]$ obtained via Fourier projection after coordinate rescaling.
Theorem 10.6.
The approximate Lyapunov exponent is computed as
$\lambda_{s}(\alpha,\beta,\sigma,k)=$ $\;[\ln(\alpha(1+\beta))-(\alpha-1)]\cdot\mathcal{F}_{[0,1]}(\widetilde{f}_{\sigma,k,s})[0]$
$+\sum_{j=1}^{k}\left(\frac{2(\alpha-1)}{j\pi}\int_{0}^{j\pi}\frac{\sin(x)}{x}dx\right)\cdot\mathcal{F}_{[0,1]}(\widetilde{f}_{\sigma,k,s})[j].$
11. Code
The code used to perform our rigorous computations is publicly available at the following link:
https://github.com/orkolorko/PlateauExperiment.jl.
This program enables the certified computation of tight intervals that contain the Lyapunov exponent of the noise-driven dynamical system described in (3), for any given values of the parameters $\alpha$, $\beta$, and $\sigma$. All computations are performed with rigorous error bounds using interval arithmetic and validated numerics.
Acknowledgements
I.N. was partially supported by the Posgraduate Program in Mathematics at UFRJ, CAPES–Finance Code 001, CNPq Projeto Universal No. 404943/2023-3, CAPES–PRINT No. 88881.311616/2018-00, and CAPES–STINT No. 88887.15574
6/2017-$0$.
C.L.V. was partially supported by Posgraduate Program in Mathematics at UFRJ, CAPES-Process No. 88882.331108/2019-01.
The authors would like to thank Prof. Yuzuru Sato for introducing them to the fascinating topic of noise-induced phenomena and for generously sharing his deep understanding of their phenomenological behavior.